1. (Exercise*) Solve the first order initial value problem y + 2y =t(ui(t) – uz(t)) subject to y(0) = 0.
Answer #1

4 +2y = t (t)-Walt) ft Yoo-o. Tabing Laplace transform of eno LH:0+ 2. Lý) = [dt) – $4,067 fly)-460) +244)= L(+44(E)) — 11+4